The RS Center takes an average of 55 days to sell its inventory and an average of 20 days to

collect payment on its sales and 30 days to pay its suppliers.

What is the inventory turnover rate?

What is the receivables turnover rate?

What is the payable turnover rate?

(If questions give you the turnover rate instead, can you calculate how long on average does it

take the firm to sell/to get payment/to pay?)
Answer #1
Ans. 1 Average days in inventory = No. of days in year / Inventory turnover
55 =   365 / Inventory turnover
Inventory turnover = 365 / 55
Inventory turnover = 6.64   times
Ans. 2 Average days in receivables = No. of days in year / Receivable turnover
20 =   365 / Receivables turnover
Receivable turnover = 365 / 20
Receivable turnover = 18.25   times
Ans. 3 Average days in payable = No. of days in year / Payable turnover
30 =   365 / Payable turnover
Payable turnover = 365 / 30
Payable turnover = 12.17   times
(Yes, we can calculate the average days in inventory, receivables and payable
by using the same formula)